OpenStudy (anonymous):

4. In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x) 1. Complete the table of values. x 100 + 23x 90(1.2x) 0 100 + 0 90(0) 1 100 + 23 90(1.2) 2 100 + 46 90(1.44) 3 100 + 69 90(1.728) 4 100 + 92 90(2.0736) 5 100 + 115 90(2.48832) 2. Use the table to determine at approximately which point the number of cells will be the same for each organism. 3. Graph the system of equations and sho

OpenStudy (anonymous):

.4. In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x) 1. Complete the table of values. x 100 + 23x 90(1.2x) 0 100 + 0 90(0) 1 100 + 23 90(1.2) 2 100 + 46 90(1.44) 3 100 + 69 90(1.728) 4 100 + 92 90(2.0736) 5 100 + 115 90(2.48832) 2. Use the table to determine at approximately which point the number of cells will be the same for each organism. 3. Graph the system of equations and show the point of intersection. 4. Explain what the points graphed for each line represent. 5. Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c. 6. Find the point(s) of intersection, and explain what the intersection represents in the context of the problem. @bibby

OpenStudy (bibby):

As usual, the intersection is the point where the 2 equations are equal to each other. that should give you 5/6. for 1, solve the different parts using a calculator and see which ones are equal

OpenStudy (anonymous):

yea i did that and its shown . 2,3,4,5,6 are the ones that i need help with.

OpenStudy (bibby):

Can you post your work? I don't want to do the arithmetic.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

overv and over again.

OpenStudy (bibby):

I mean stuff like this: 100 + 115 90(2.48832) I don't know if the two numbers are equal

OpenStudy (anonymous):

oh 1.100(first) and 0(other) 2. 123 and 108 3. 146 and 126.6 4. 169 and 155.52 5. 192 and 186.624 6. 215 and 223.9488

OpenStudy (bibby):

ok so 2 wants us to approximate where the 2 equations will intersect. Do you have any ideas?

OpenStudy (anonymous):

well maybe... well from this it doesnt look like they will meet...

OpenStudy (bibby):

if you look closer, they're getting closer as you get to 6. It starts off at 15 units, gets to 8 between 5 and 6

OpenStudy (anonymous):

yea i didnt see that so hold on.

OpenStudy (anonymous):

238 and 268.73856

OpenStudy (bibby):

Are those the next numbers?

OpenStudy (anonymous):

yes

OpenStudy (bibby):

Yeah that confirms it. The intersection happens between 5 and 6. maybe 5.5 or so

OpenStudy (anonymous):

ok so ill put 5.5